The hypotenuse of a right triangle measures 10 cm. one of the other two sides is 2 m longer than the second. Find the unknown side lengths.
Answer by richwmiller(17219)

a^2+b^2=100,
a=b+2
The 10 should make you think of the [3,4,5] right triangle since 10 is a multiple of 5.
so the other s sides will be 2*3=6 and 2*4=8